OpenStudy (anonymous):

Please help!!!!!!!!!!!!!!!!!!!!!!!!!!! int_{1}^{8}4x ^{1/3}dx

OpenStudy (anonymous):

int_{1}^{8}4x ^{1/3}dx

OpenStudy (anonymous):

\[\int\limits_{1}^{8}4x ^{1/3}dx\]

OpenStudy (anonymous):

You'll want to use the identity \(\int x^r dx = \frac{1}{r+1}x^{r+1} + C \text{ if } r \neq -1\). In your case, r = 1/3.

OpenStudy (anonymous):

Maybe indentity is the wrong word but you get the idea.

OpenStudy (rock_mit182):

\[\int\limits_{a}^{b} k*x ^{r} dx = k \int\limits_{a}^{b}x ^{r}\] where k is constant and r is the exponent

OpenStudy (rock_mit182):

onde k é uma consante e r é o exponente da funcao

OpenStudy (anonymous):

certo

OpenStudy (rock_mit182):

sim meu amigo, voce estudia engenheria ?

OpenStudy (anonymous):

sim amigo, estou vendo essa matéria agora...

OpenStudy (anonymous):

5*3^2 = 45

OpenStudy (anonymous):

está certa a resposta?

OpenStudy (rock_mit182):

I'm not allowed to give you a final answer, those are the rules here

OpenStudy (rock_mit182):

amigo, isso é um oito ?

OpenStudy (rock_mit182):

o é uma S ?

OpenStudy (anonymous):

found this result, now I wonder if it is correct.

OpenStudy (rock_mit182):

i wonder if the top of the integral is a number or a letter

OpenStudy (rock_mit182):

I mean, cannot see it clear, is an 8 or S ?

OpenStudy (anonymous):

is the number 8

OpenStudy (rock_mit182):

\[= \frac{ x ^{\frac{ 1 }{ 3 }+1} }{ 1/3+1 }\] evaluated in 0 and 8 (i guess,ñao olho bem a expressao arriba na integral )

OpenStudy (anonymous):

1 and 8 ?

OpenStudy (anonymous):

int_{1}^{8}4x ^{1/3}dx

OpenStudy (anonymous):

?

OpenStudy (rock_mit182):

\[\int\limits_{0}^{8}4x ^{\frac{ 1 }{ 3 }}dx=4\int\limits_{0}^{8}x ^{\frac{ 1 }{ 3 }}dx=4[\frac{ x ^{\frac{ 1 }{ 3 }+1} }{ \frac{ 1 }{ 3 }+1 }]_{0}^{8}\]

OpenStudy (rock_mit182):

\[4[\frac{ 3\sqrt[3]{8^{4}} }{ 4 }]\]=